# A Tale of Single-channel Electroencephalogram: Devices, Datasets, Signal Processing, Applications, and Future Directions

### v1

@@ VI-B Emotion Recognition
Emotion is a multifaceted psychological and physiological state observed in humans and other organisms in response to external or internal stimuli. It significantly influences behavior, cognition, adaptation to the environment, decision-making, and social interactions. Emotions entail internal physiological changes including variations in heart rate, respiration, and brain electrical activity [117]. Arousal level denotes the extent of physiological and psychological activation experienced during an emotion, ranging from negative to positive states [118]. Valence refers to the pleasantness of an emotion, indicating whether it is positive or negative and is also measured on a negative to positive scale [119]. Integrating arousal level and valence allows for a comprehensive classification of emotional states [120], as illustrated in Figure 5, providing a clear framework for emotion recognition and regulation.
@@
With the advent of wireless head-mounted single-channel EEG devices, the requirement for conductive glue has diminished, enhancing their portability and flexibility. This development has positioned them as the primary option for emotion recognition research, owing to their ease of use and mobility [121]. Table IV presents various applications of single-channel EEG in emotion recognition.
Zhao et al. introduced an interpretable method for emotion classification using single-channel EEG signals based on the Emo_Food dataset [41]. They utilized wavelet packet decomposition and statistical feature extraction to extract multi-domain features from EEG signals, employing extreme gradient boosting (XGBoost) for classification and achieving global and local interpretations of emotion classification using the Shapley additive explanation (SHAP) method [122]. Yan et al. proposed an emotion recognition method using single-channel EEG signals from DEAP and DREAMER datasets. They employed topological nonlinear dynamic analysis, including phase space reconstruction and persistent homology, to study EEG characteristics during emotional states [123]. Khan et al. investigated emotion induction through video games using single-channel EEG equipment. Participants across various age groups played different games to induce emotions, and EEG signals were analyzed alongside sentiment classification using machine learning techniques [124]. Cao et al. developed an emotion recognition method tailored for single-channel EEG signals. Their approach involved WT for signal decomposition, extraction of time-frequency information from multiple frequency bands, and sliding time windows for feature calculation and emotion classification [125]. Jalilifard et al. studied the classification of emotional states using single-channel EEG, focusing on the left prefrontal cortex (Fp1 location). They utilized emotional stimulus videos and applied K-nearest neighbor and SVM algorithms for successful classification of fear and relaxation emotions [81]. Acharya et al. proposed a genetic programming (GP) framework with a new fitness function ”Gap Score (G Score)” for imbalanced emotion recognition datasets. They used data collected from the NeuroSky MindWave II single-channel EEG device to classify four emotion types: happy, scary, sad, and neutral [126]. Li et al. proposed a technique based on brain rhythm sequencing (BRS) for efficient emotion recognition from EEG signals and achieved single-channel selection through dynamic time warping. The results showed that using single-channel data and 10-second brain rhythm sequences, a classification accuracy of 70-82% was achieved in the emotion recognition task [127].
@@
VI-C Educational Research
Single-channel EEG can assess cognitive load during learning tasks and stages by monitoring changes in brain activity, aiming at optimizing teaching content, methods and outcomes [13]. In a classroom environment, single-channel EEG can monitor changes in attention to teaching content, helping teachers adjust strategies and rhythm to improve outcomes. When evaluating educational technology and online learning tools, single-channel EEG provides objective evaluation metrics by measuring changes in brain activity with different tools.
Wei and Ma utilized the NeuroSky MindBand[34] to investigate the effects of visual attention and reading duration on children and adults while reading printed picture books [128]. Chen and Lin employed the NeuroSky MindSet headset in mobile reading environments to assess changes in learners’ attention across sitting, standing and walking scenarios [129]. Shadiev et al. investigated the impact of speech-to-text applications on students’ academic performance, attention and meditation [130] using NeuroSky MindWave headsets [33]. Ma and Wei employed the NeuroSky MindBand to study children’s attention during reading across varied media formats (traditional, pop-up, audio, e-books), focusing on grades 3-6 and analyzing gender and grade-level influences [131]. Chen and Wu used NeuroSky MindSet headset to examined how different video lecture methods (lecture screenshots, picture-in-picture, voice-over) affect students’ sustained attention [132]. Similarly, Wang and Hsu applied NeuroSky MindSet headset to measure the changes in learners’ attention state under the same learning content with three different text display types (static, dynamic, mixed text display types) [28].
In traditional teacher-student interaction, Huang et al. proposed to stimulate students’ positive emotions by rewarding correct answers with interesting images, and found that such incentives are effective [133], using the NeuroSky MindWave Mobile headset to measure attention levels. In the realm of intelligent tutoring systems, [134] conducted a comprehensive study comparing the effects of clicker technology versus mobile voting technology on students’ anxiety, self-efficacy, engagement, academic performance and attention, employing the NeuroSky MindSet headset to measure attention levels. Wang et al. measured learners’ attention levels in computer-based teaching and learning across different difficulty levels using the NeuroSky MindSet headset [28]. Similarly, [135, 136, 137] also utilized the NeuroSky MindSet headset to gauge students’ attention levels in online synchronous teaching and learning activities. Wong et al. investigated participants’ attention levels in motor skill acquisition using the NeuroSky MindWave Mobile headset, employing a computerized visual motor task akin to mirror drawing [138]. Lin et al. conducted brainwave concentration and relaxation training in an emotional learning system with the NeuroSky MindWave Mobile, automatically halting teaching materials and recording learning times upon unexpected negative emotions [29].
@@ VI-C Educational Research
TABLE IV: Article summary of Emotion recognition based on Single-channel EEG Systems
@@
  Ref Device Datasets Selected Channel Features Model Result(%)
@@
  [122] Neurosky Mindwave Mobile II Emo_Food Dataset Fp1 Four frequency bands($\alpha$,low $\beta$,high $\beta$, $\gamma$) XGBoost 92.31
@@
[127] Multi-channel system device MER Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS 72.11
@@
SEED Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS 72.40
@@
DEAP Dataset C3 $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS 74.51
@@
MAHNOB Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS 82.18
@@
[124] Neurosky Mindwave Mobile II Self-built Dataset Fp1 Time, Frequency and Time-frequency Domains Boosted Trees classifier 82.26
@@
[125] Neurosky Mindwave Emo_Food Dataset Fp1 Four frequency bands($\alpha$,low $\beta$, high $\beta$, $\gamma$) TF-LSTM 98.36
@@
[81] Neurosky Mindwave Mobile Self-built Dataset Fp1 Four frequency bands($\theta$, $\alpha$, $\beta$, $\gamma$) KNN 94.21
@@
[126] Neurosky Mindwave Mobile II Self-built Dataset Fp1 Attention, Meditation, $\delta$, $\theta$, Low $\alpha$, High $\alpha$, Low $\beta$, High $\beta$ , Low $\gamma$, and High $\gamma$ GP 87.61
VI-D Clinical Diagnosis and Treatment Assessment
@@
The significance of single-channel EEG in medical diagnosis has grown rapidly due to its ability to provide swift and effective brain state assessments, crucial in acute medical scenarios such as stroke onset or epileptic seizures, where rapid decision-making is vital. In mental health, where conditions like depression exhibit detectable EEG pattern changes, and single-channel EEG supports ongoing monitoring essential for clinical diagnosis and treatment assessment. In neurodegenerative conditions like Alzheimer’s Disease (AD) [139] and Mild Cognitive Impairment (MCI) [140], single-channel EEG detects brain wave anomalies indicative of cognitive decline, enabling early detection and intervention. Its simplicity and accessibility position single-channel EEG as a promising tool for large-scale screening.
VI-D1 Depression
@@
Single-channel EEG offers a noteworthy advantage due to its simplicity and rapid deployment capability, facilitating timely and potentially more frequent assessments in mental health contexts. Depression, a prevalent and debilitating disorder, necessitates diagnostic tools that are effective, accessible, and efficient. Studies using single-channel EEG for diagnosing depression have generally focused on analyzing brain wave patterns and connectivity measures to distinguish between depressed patients and healthy controls [141, 79]. Much of this research aims to identify specific EEG biomarkers indicative of depression. For instance, alterations in frontal $\alpha$ asymmetry have been commonly observed in depressive patients, suggesting a potential diagnostic signature detectable through single-channel EEG [142]. A pivotal study showed that single-channel EEG, using standard ML techniques to analyze signals from a single frontal electrode, effectively identifies depression by detecting reduced left frontal activity. This approach offers diagnostic accuracy comparable to multi-channel systems [79]. Another method involves inducing cognitive or emotional stimuli and recording EEG responses to tasks that elicit brain activity related to depression. Analyzing these differences in electrical brain responses between depressed patients and controls provides insights into the neural mechanisms of emotional dysregulation in depression [141]. Various EEG features studied in depression include spectral power in different frequency bands ($\alpha$, $\beta$, $\gamma$, $\delta$, $\theta$), asymmetry indices, and coherence measures. The alpha asymmetry index, extensively investigated as a potential biomarker for depression, assesses the disparity in alpha wave activity between the brain’s left and right hemispheres. Greater left-sided inactivity is often associated with depressive symptoms [143].
@@
Chung et al. proposed a method for assessing depression using a single-channel EEG device (NeuroSky Mindwave Mobile II) and a DNN model, demonstrating efficient assessment through multi-layer ensemble learning technology [79]. Zhu et al. introduced a method for detecting depression through single-channel EEG and graph-based techniques in 2022, achieving highest accuracy of 92.0% with the T4 channel during eyes-closed states [144]. Shen et al. presented an enhanced channel selection approach using kernel target alignment for depression detection in multi-channel EEG, showing improved classification performance and satisfactory results with a single channel aligned with cortical activity patterns associated with depression [145]. Wan et al. explored a ML method for distinguishing major depressive disorder (MDD) from normal controls using single-channel EEG data from forehead locations (Fp1 and Fp2 electrodes). Their results indicated superior classification performance at Fp1, showcasing single-channel EEG’s capability comparable to multi-channel EEG in MDD differentiation [146]. Ay et al. proposed a deep hybrid model combining CNN and LSTM architecture to detect depression through EEG signals, achieving high classification accuracy of EEG signals in the right and left hemispheres 99.12% and 97.66%, respectively [83]. Additional studies [147, 148, 149] underscored the effectiveness of single-channel EEG analysis across various methods for detecting depression, emphasizing its feasibility and efficacy in linear and nonlinear EEG signal analysis techniques.
VI-D2 AD & MCI
@@ VI-D3 Epilepsy
Epilepsy, characterized by recurrent and unprovoked seizures, is one of the most prevalent neurological disorders globally [6]. Traditionally, multi-channel EEG has been the standard diagnostic tool for clinicians. However, its complexity, high cost, and need for specialized equipment and expertise have limited its widespread use, particularly in resource-limited settings. In contrast, single-channel EEG offers a simpler, cost-effective alternative with significant clinical utility. Dweiri et al. created a new ML algorithm using XGBoost for seizure detection in wearable systems, classifying single-channel EEG from the CHB-MIT database, which demonstrated high sensitivity of 89% for seizure detection [152]. Another notable study placed sensors in varying numbers of channels to detect epilepsy separately using the CHB-MIT dataset. The results indicated that the single-channel method’s performance in identifying epileptic seizures in continuous EEG recordings was comparable to the multi-channel method, with high sensitivity and a low false alarm rate [153]. Lu et al. compared the diagnostic efficacy of single-channel and standard multi-channel EEG in a clinical trial involving patients suspected of nocturnal seizures. Results suggested that while single-channel EEG is less detailed, it still correctly identified significant seizure activities in the majority of cases [154]. This finding supports the potential of single-channel EEG as a screening tool, particularly in environments where access to full EEG setups is not feasible.
VI-E Other Applications

### v2

@@ VI-B Emotion Recognition
Emotion is a multifaceted psychological and physiological state observed in humans and other organisms in response to external or internal stimuli. Emotions entail internal physiological changes including variations in heart rate, respiration, and brain electrical activity [11]. Arousal level denotes the extent of physiological and psychological activation experienced during an emotion, ranging from negative to positive states [139]. Valence refers to the pleasantness of an emotion, indicating whether it is positive or negative and is also measured on a negative to positive scale [140]. Integrating arousal level and valence allows for a comprehensive classification of emotional states [141], as illustrated in Fig. 5, providing a clear framework for emotion recognition and regulation.
@@
With the advent of wireless head-mounted single-channel EEG devices, the requirement for conductive glue has diminished, enhancing their portability and flexibility. This development has positioned them as the primary option for emotion recognition research, owing to their ease of use and mobility [142]. Table V presents various applications of single-channel EEG in emotion recognition. Zhao et al. introduced an interpretable method for emotion classification using single-channel EEG signals based on the Emo_Food dataset [62]. They employed wavelet packet decomposition and statistical feature extraction for multi-domain features, using extreme gradient boosting (XGBoost) [143] and Shapley additive explanation (SHAP) method for interpretable emotion classification [62]. Yan et al. proposed an emotion recognition method using single-channel EEG signals from DEAP [144] and DREAMER datasets [145]. They employed topological nonlinear dynamic analysis, including phase space reconstruction and persistent homology, to study EEG characteristics during emotional states [146]. Khan et al. investigated emotion induction through video games using single-channel EEG equipment. Participants across various age groups played different games to induce emotions, and EEG signals were analyzed alongside sentiment classification using ML techniques [147]. Cao et al. developed an emotion recognition method tailored for single-channel EEG signals. Their approach involved WT for signal decomposition, extraction of time-frequency information from multiple frequency bands, and sliding time windows for feature calculation and emotion classification [148]. Jalilifard et al. studied the classification of emotional states using single-channel EEG, focusing on the left prefrontal cortex (Fp1 location). They utilized emotional stimulus videos and applied KNN and SVM algorithms for successful classification of fear and relaxation emotions [103]. Acharya et al. proposed a genetic programming (GP) framework with a new fitness function “Gap Score (G Score)” for imbalanced emotion recognition datasets. They used data collected from the NeuroSky MindWave II single-channel EEG device to classify four emotion types: happy, scary, sad, and neutral [149]. Li et al. proposed a brain rhythm sequencing (BRS) technique with dynamic time warping for single-channel selection, achieving 70-82% classification accuracy using 10-second brain rhythm sequences for emotion recognition [150].
@@
VI-C Neurofeedback
@@
TABLE IV: Article summary of Neurofeedback based on Single-channel EEG Systems
  Ref Device Selected Channel Frequency Band Subject Number Application Domain
  [151] NeuroSky MindWave Mobile Fp1 $\alpha$ 16 Anxiety
[152] Neurosky EEG headset Fp1 high $\alpha$, low $\alpha$, $\beta$, low $\beta$, $\theta$ 40 Enchancing cognitive performance
[153] Self-built Device Fpz $\theta$, $\alpha$, and $\beta$ 6 Home nursing
Neurofeedback is a safe, non-invasive technique grounded in learning theories from neuroscience and behavioral science, aimed at enhancing brain function and structure. By measuring neural activity associated with specific functions and delivering real-time feedback through visual, auditory, or tactile stimuli, neurofeedback encourages individuals to employ psychological strategies to voluntarily modulate targeted neural activity, thereby enhancing behavior and cognitive function. Neurofeedback has demonstrated positive outcomes in managing depression and anxiety [154]. However, limited training sessions, small sample sizes, lack of standardization and unclear long-term effects constrain its efficacy, requiring further research to address these limitations [155].
A comprehensive investigation with NeuroSky’s wireless fronto-polar EEG monitoring revealed significant inverse correlation between $\alpha$ wave activity and anxiety levels ($p$leq 0.029) in adolescent athletes during eyes-closed conditions, demonstrating single-channel EEG systems’ efficacy for psychological assessment and athletic anxiety management [151]. A cognitive intervention study utilizing cost-effective single-channel EEG devices [152] implemented five brain-training protocols across 20 sessions in five Thai medical centers, demonstrating statistically significant enhancements ($p$leq 0.05) in visual memory, attention, and recognition among cognitively-screened elderly participants, thus establishing single-channel EEG neurofeedback as a viable therapeutic intervention for cognitive enhancement in aging populations. Sandra et al. [156] elucidate the clinical applications of EEG in attention deficit hyperactivity disorder (ADHD) research and treatment, emphasizing EEG-based neurofeedback therapeutic efficacy while acknowledging diagnostic heterogeneity. Hande et al.’s [157] investigation implemented neurofeedback training protocols targeting $\theta$ and sensorimotor rhythm (12-15 Hz) bands in adult cohorts, evaluating selective attention through stroop task measurements.
Extant neurofeedback research faces methodological limitations from insufficient training sessions and suboptimal task complexity, requiring enhanced protocols with increased sessions and challenging paradigms for conclusive clinical efficacy evidence, as summarized in Table IV for single-channel EEG interventions.
VI-D Educational Research
Single-channel EEG monitors brain activity changes to assess cognitive load during learning tasks and stages, enabling optimization of teaching content, methods and outcomes [13]. In classroom environments, it monitors attention changes to teaching content, helping teachers improve strategies and rhythm. When evaluating educational technology and online learning tools, single-channel EEG provides objective evaluation metrics by measuring changes in brain activity with different tools.
Research examining single-channel EEG in educational contexts has yielded diverse findings. Initial investigations by Wei et al. utilized NeuroSky MindBand[47] to study visual attention and reading duration effects on children and adults during picture book reading [158]. Subsequent mobile environment research assessed learner attention variations across sitting, standing and walking scenarios through NeuroSky MindSet headset [159]. Further studies investigated speech-to-text applications’ impact on academic performance, attention and meditation through NeuroSky MindWave [160, 45]. Additional research explored children’s attention during reading across varied media formats including traditional, pop-up, audio and e-books, focusing on grades 3-6 while analyzing gender and grade-level influences [161]. Research examining video lecture methods analyzed how different presentation styles affected student sustained attention [162], while parallel studies investigated attention state changes under varying text display types [39].
Studies in traditional classroom settings demonstrated that rewarding correct answers with interesting images effectively stimulated positive student emotions [55]. Comprehensive research in intelligent tutoring systems compared clicker technology with mobile voting technology, examining effects on student anxiety, self-efficacy, engagement and academic performance [163]. Additional investigations measured attention levels across varying difficulty levels in computer-based learning [39] and online synchronous teaching activities [164, 165, 166]. Research in motor skill acquisition investigated attention during computerized visual motor tasks [167], while studies in emotional learning systems examined brainwave concentration and relaxation training, incorporating automatic material adjustment based on emotional states [40].
TABLE V: Article summary of Emotion recognition based on Single-channel EEG Systems
@@ VI-D Educational Research
  Ref Device Datasets Selected Channel Features Model Emotion Classes Emotion Result(%)
@@
  [62] Neurosky Mindwave Mobile II Emo_Food Dataset Fp1 Four frequency bands($\alpha$,low $\beta$,high $\beta$, $\gamma$) XGBoost 3 Neutral, Positive, Negative 92.31
@@
[150] Multi-channel system device MER Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS Continuous 5 emotional factors: familiarity, liking, understanding, arousal, and valence(1-9) 72.11
@@
SEED Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS 3 Neutral, Positive, Negative 72.40
@@
DEAP Dataset C3 $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS Continuous 2 emotional factors: arousal, valence(1-9) 74.51
@@
MAHNOB Dataset $\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$ BRS Continuous 2 emotional factors: arousal, valence(1-9) 82.18
@@
[147] Neurosky Mindwave Mobile II Self-built Dataset Fp1 Time, Frequency and Time-frequency Domains Boosted Trees classifier 4 Happy, Bored, Relaxed, Stressed 82.26
@@
[148] Neurosky Mindwave Emo_Food Dataset Fp1 Four frequency bands($\alpha$,low $\beta$,high $\beta$, $\gamma$) TF-LSTM 3 Neutral, Positive, Negative 98.36
@@
[103] Neurosky Mindwave Mobile Self-built Dataset Fp1 Four frequency bands($\theta$, $\alpha$, $\beta$, $\gamma$) KNN 3 Neutral, Relaxation, Scary 94.21
@@
[149] Neurosky Mindwave Mobile II Self-built Dataset Fp1 Attention, Meditation, $\delta$, $\theta$, Low $\alpha$, High $\alpha$, Low $\beta$, High $\beta$, Low $\gamma$, and High $\gamma$ GP 4 Happy, Sad, Horror, Neutral 87.61
  • XGBoost: Extreme Gradient Boosting. BRS: Brain Rhythm Sequences. TF-LSTM: Time-Frequency Long Short-Term Memory. KNN: K-Nearest Neighbors. GP: Genetic Programming. Continuous: measured on a continuous scale, not discrete.
VI-E Clinical Diagnosis and Treatment Assessment
@@
Single-channel EEG has gained significance in medical diagnosis through swift and effective brain state assessments. It is particularly crucial for rapid decision-making in acute conditions like stroke onset and epileptic seizures. In mental health conditions, single-channel EEG enables continuous monitoring through detectable pattern changes, supporting clinical diagnosis and treatment evaluation. For neurodegenerative conditions like Alzheimer’s Disease (AD) [168] and Mild Cognitive Impairment (MCI) [169], single-channel EEG detects brain wave anomalies indicative of cognitive decline, enabling early detection and intervention. Its simplicity and accessibility position single-channel EEG as a promising tool for large-scale screening.
VI-E1 Depression
@@
Single-channel EEG offers a noteworthy advantage due to its simplicity and rapid deployment capability, facilitating timely and potentially more frequent assessments in mental health contexts. Depression, a prevalent and debilitating disorder, necessitates diagnostic tools that are effective, accessible, and efficient. Studies using single-channel EEG for diagnosing depression have generally focused on analyzing brain wave patterns and connectivity measures to distinguish between depressed patients and healthy controls [170, 101]. Much of this research aims to identify specific EEG biomarkers indicative of depression. For instance, alterations in frontal $\alpha$ asymmetry have been commonly observed in depressive patients, suggesting a potential diagnostic signature detectable through single-channel EEG [171]. A pivotal study demonstrated that single-channel EEG analysis of frontal electrode signals through ML techniques effectively identifies depression via reduced left frontal activity, achieving diagnostic accuracy comparable to multi-channel systems [101]. Another approach examines EEG responses to cognitive or emotional stimuli tasks, analyzing electrical brain response differences between depressed patients and controls to reveal neural mechanisms underlying emotional dysregulation [170]. Depression studies investigate various EEG features including spectral power across frequency bands ($\alpha$, $\beta$, $\gamma$, $\delta$, $\theta$), asymmetry indices, and coherence measures, with particular emphasis on the alpha asymmetry index as a potential depression biomarker through assessment of alpha wave activity disparity between hemispheres, where greater left-sided inactivity correlates with depressive symptoms [172].
@@
Chung et al. proposed a method for assessing depression using a single-channel EEG device (NeuroSky Mindwave Mobile II), demonstrating efficient assessment through multi-layer ensemble learning technology [101]. Zhu et al. introduced a method for detecting depression through single-channel EEG and graph-based techniques in 2022, achieving highest accuracy of 92.0% with the T4 channel during eyes-closed states [173]. Shen et al. presented an enhanced channel selection approach using kernel target alignment for depression detection in multi-channel EEG, showing improved classification performance and satisfactory results with a single channel aligned with cortical activity patterns associated with depression [174]. Wan et al. explored a ML method for distinguishing major depressive disorder (MDD) from normal controls using single-channel EEG data from forehead locations (Fp1 and Fp2 electrodes). Their results indicated superior classification performance at Fp1, showcasing single-channel EEG’s capability comparable to multi-channel EEG in MDD differentiation [175]. Ay et al. proposed a deep hybrid model combining CNN and LSTM architecture to detect depression through EEG signals, achieving high classification accuracy of EEG signals in the right and left hemispheres 99.12% and 97.66%, respectively [105]. Additional studies [176, 177, 178] underscored the effectiveness of single-channel EEG analysis across various methods for detecting depression, emphasizing its feasibility and efficacy in linear and nonlinear EEG signal analysis techniques.
VI-E2 AD & MCI
@@ VI-E3 Epilepsy
Epilepsy, characterized by recurrent and unprovoked seizures, represents one of the most prevalent neurological disorders globally [6]. While multi-channel EEG remains the standard diagnostic tool, its complexity, high cost, and specialized requirements limit widespread application, particularly in resource-limited settings. Single-channel EEG emerges as a simpler, cost-effective alternative with significant clinical utility. Dweiri et al. created a new ML algorithm using XGBoost for seizure detection in wearable systems, classifying single-channel EEG from the CHB-MIT database, which demonstrated high sensitivity of 89% for seizure detection [181]. Another notable study placed sensors in varying numbers of channels to detect epilepsy separately using the CHB-MIT dataset. The results indicated that the single-channel method’s performance in identifying epileptic seizures in continuous EEG recordings was comparable to the multi-channel method, with high sensitivity and a low false alarm rate [182]. Lu et al. compared the diagnostic efficacy of single-channel and standard multi-channel EEG in a clinical trial involving patients suspected of nocturnal seizures. Results suggested that while single-channel EEG is less detailed, it still correctly identified significant seizure activities in the majority of cases [183]. This finding supports the potential of single-channel EEG as a screening tool, particularly in environments where access to full EEG setups is not feasible.
VI-F Other Applications